Harrelson Inc. currently has $750,000 in accounts receivable, and its days sales outstanding (DSO) is 55 days. It wants to reduce its DSO to 35 days by pressuring more of its customers to pay their bills on time. If this policy is adopted the company's average sales will fall by 15 percent. What will be the level of accounts receivable following the change? Assume a 365 day year.
Answer #2

GIVEN:

Accounts Receivable                      =          $750,000

Days Sales Outstanding (DSO)     =          55 days

Days Sales Outstanding new (DSO new)           =          35 days

Sales will fall by                                                        =          15%

 

 

 

 

 

 


Accounts Receivable = (365/55) x Sales

Sales A          = Accounts Receivable A  x  (365/DSO A) =
                       = $750,000 x (365/55) = $4,977,272.73

Sales A          = $4,977,273


After the new policy has been implemented…

DSO new                   = 35 days
Sales new                 = (100% - 15%) of Sales A =
                                   = 0.85 x $4,977,272.73 = $4,230,681.82

Sales new                 = $4,230,682

Now,

 

Accounts Receivable new  =          (DSO1/365) x Sales1 =
                                           =         (35/365) x $4,230,681.82 = $405,681.818

Accounts Receivable new  =          $405,682


answered by: yEN
Answer #1
Currently,
DSO0 = 55 days
Note that
Accounts Receivable = (DSO/365) x Sales
so
Sales0 = Accounts Receivable0 x (365/DSO0) =
= $750,000 x (365/55) = $4,909,090.91
After the new policy has been implemented…
DSO1 = 35 days
Sales1 = (100% - 15%) of Sales0 =
= 0.85 x $4,909,090.91 = $4,172,727.27
Now,
Accounts Receivable1 = (DSO1/365) x Sales1 =
= (35/365) x $4,172,727.27 = $400,124.53